A company makes an HD-TV that has a 63.4 inch diagonal. If the width of the screen is approximately 51.8 inches, what is the approximate height of the screen? Round your answer to the nearest Tenth.
Answer by oscargut(686)

D=diagonal
W=width
H=height
then using Phytagoras theorem D^2=W^2+H^2
then (63.4)^2=(51.8)^2+H^2
then 4019.56=2683.24+H^2
then H^2=1336.32
then H=sqrt(1336.32)=36.557=36.6(aprox)
Answer: height is 36.6 inches